OpenStudy (anonymous):

write an equation of the line that passes through the given point and has the given slope m. (-3,-11) m=1/2 write an equation of the line that passes through the given point and has the given slope m. (-3,-11) m=1/2 @Mathematics

OpenStudy (saifoo.khan):

\[y = \frac 12x-\frac{19}{2}\]

OpenStudy (anonymous):

how did u do that

OpenStudy (saifoo.khan):

Magic?

OpenStudy (anonymous):

no i mean the steps

OpenStudy (saifoo.khan):

lol. what are your teachers going? just wondering..

OpenStudy (anonymous):

well most of them dont care and just give us work that isnt tought . i go on here because sometimes people help me. thats why im thinking of droping out of algebra 1 to go to advanced math since the teacher doesnt teach

OpenStudy (saifoo.khan):

teachers suck there!

OpenStudy (anonymous):

yes they do alot i just moved here and i wish i could go back

OpenStudy (saifoo.khan):

ohh.. dont u just need the answers?

OpenStudy (anonymous):

ya can u help me with a few

OpenStudy (anonymous):

because i think ur a big help and everytime i go on i look for u to help me

OpenStudy (saifoo.khan):

how many do u have?

OpenStudy (anonymous):

6 -32 even plus she added 40,44,and 50 thats alot but not all r the same thing

OpenStudy (saifoo.khan):

u want direct answers right?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

post them! xD

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

write an equation of the line that passes through the given points (3,2) (4,9)

OpenStudy (anonymous):

write an equation of the line that passes through the given points (-2,8)( -6,0)

OpenStudy (saifoo.khan):

(3,2) (4,9) y = 7x - 19

OpenStudy (saifoo.khan):

(-2,8)( -6,0) y = 2x + 12

OpenStudy (anonymous):

write an equation of the line that passes through the given points (-5, 3/4) (-2,-3/4)

OpenStudy (saifoo.khan):

\[y = -\frac12x - \frac74\]

OpenStudy (anonymous):

and this one has a graph which is plotting two points.( 3,2) and (-2,0) its says to write an equation of the line shown. the line goes across the y axis

OpenStudy (anonymous):

and this one has a graph which is plotting two points.(-1,3) and (2,-0.5) its says to write an equation of the line shown.

OpenStudy (saifoo.khan):

( 3,2) and (-2,0) \[y = \frac 25x +\frac45\]

OpenStudy (anonymous):

k

OpenStudy (saifoo.khan):

.(-1,3) and (2,-0.5) \[y = -\frac76x +\frac{11}{6}\]

OpenStudy (anonymous):

its the same probelm again (-3,2) (-2,-1)

OpenStudy (anonymous):

do the same thing

OpenStudy (saifoo.khan):

y = -3x-7

OpenStudy (anonymous):

write an equation for a linear equation f that has a given value f(-2)=-2,f(4)= -8

OpenStudy (anonymous):

write an equation for a linear equation f that has a given value f(-4)=9,f(-8)= -11

OpenStudy (karatechopper):

hey saif guess what?! i learned tht just 2 weeks ago! i feel to b a master at it now:)

OpenStudy (saifoo.khan):

what u learned? @kfc.

OpenStudy (saifoo.khan):

@wlb, i remember u said u had 6. :o

OpenStudy (karatechopper):

the question tht wlbrt posted yanoo? i love those

OpenStudy (saifoo.khan):

Ohh.. Great.

OpenStudy (saifoo.khan):

So Master Madam. start helping her! =D

OpenStudy (karatechopper):

she gots to give me a prob!

OpenStudy (saifoo.khan):

didnt she already?

OpenStudy (anonymous):

write an equation for a linear equation f that has a given value f(-4)=9,f(-8)= -11

OpenStudy (anonymous):

write an equation for a linear equation f that has a given value f(-2)=-2,f(4)= -8

OpenStudy (anonymous):

i need help

OpenStudy (saifoo.khan):

rest for tmmrw.

OpenStudy (karatechopper):

now tht one i dont kno cuz i can only do some basics like the one u posted

OpenStudy (anonymous):

ohhh

OpenStudy (saifoo.khan):

im tired.

OpenStudy (karatechopper):

saif! what did i tell u to do!

OpenStudy (karatechopper):

we have already talked about this.

OpenStudy (saifoo.khan):

Oh Yea Kfc, now i remember.. but i have no more energy. literally.

OpenStudy (karatechopper):

energy drink? i am not saying to drink monster but u could atleast drink gatorade

OpenStudy (karatechopper):

r u in the US right now?

OpenStudy (anonymous):

its alright ill guess ill do tomorr ow thanks for the help

OpenStudy (saifoo.khan):

lol, im not in US. :p that wont help, im awake from like 20hrs.

OpenStudy (karatechopper):

omg

OpenStudy (saifoo.khan):

hmm. =(